# STACKFEED: Structured Textual Actor-Critic Knowledge base editing with FEEDback

### v1

@@ 4.2 Knowledge Base Editing as State Search
Given a query $q_{i}$ and a set of retrieved documents $\Gamma(q_{i},\mathcal{K})$, the LLM $\mathcal{B}$ generates an answer $o_{i}$. When errors arise due to incomplete or incorrect information in the retrieved documents, our goal is to identify the optimal configuration of $\mathcal{K}$ that improves the accuracy of the system’s responses. Thus, we define our state search problem as finding the best state $s^{*}$ of the KB.
@@
Environment: We model the environment simply as a ‘‘patch’’ function, that takes the diff generated by the agent and patches the KB to produce the new state.
@@
$s^{*}=\arg\max_{s\in\mathcal{S}}R(s)=\arg\max_{s\in\mathcal{S}}\frac{1}{|T|}\sum_{(q_{i},a_{i},c_{i},f_{i})\in T}g(\mathcal{B}(q_{i},\Gamma(q_{i},s)),c_{i})$ (2)
@@
The reward function $R(s)$ is derived from the expert feedback on the system’s generated answers and captures improvements in terms of correctness, coherence, and completeness of the information in the KB. By optimizing for $s^{*}$, we ensure that the final state of the KB maximizes the overall accuracy and effectiveness of the RAG system, rather than focusing on an intermediate sequence of state transitions.
@@
Monte Carlo Tree Search: We employ Monte Carlo Tree Search (MCTS) similar to PromptAgent  (Wang et al., 2023) to search for the optimal state $s^{*}$. However, this introduces several challenges: (1) The search space for all possible KB edits is vastly larger than that of standard prompt edits typically explored in the literature (Pryzant et al., 2023; Wang et al., 2023; Juneja et al., 2024; Gupta et al., 2024), making exhaustive search infeasible. (2) Generating actions and subsequent states, as done in methods like PromptAgent , is difficult in the KB editing context since fitting the entire KB into the prompt of a language model is impractical. Despite advancements in handling long contexts (Wang et al., 2020; Kitaev et al., 2020; Press et al., 2022; Su et al., 2024b), these models often struggle to leverage extensive contexts effectively Liu et al. (2024). (3) Finally, the LM would need to output the entire edited KB, which is challenging due to the inherent difficulty LMs face in generating long, coherent outputs (Bai et al., 2024).
To address these challenges, we decouple the KB edits by isolating document-level modifications based on the required updates. Since individual documents can be large, we further break down the edits into manageable sections, enabling a structured editing mechanism that focuses on specific portions of a document at a time. In the next section, we introduce STACKFEED, an agent designed to efficiently perform these structured edits based on feedback.
4.3 STACKFEED
Figure 2: STACKFEED  Multi-actor, centralized critic architecture: On receiving a reward from the environment, the critic generates a reflection over the failures to calculate the textual gradient $\nabla$. The critic uses this reflection to select the documents responsible for the error and proceeds to assigns credit to the actors in the form of document-wise reflections. The actors then proceed to iteratively edit the documents. All the document-wise edits are then pooled to define the KB edit.
The proposed approach STACKFEED is designed to enhance a RAG system by refining the underlying Knowledge Base ($\mathcal{K}$) using expert feedback. Our approach employs a multi-actor, centralized critic architecture, where each actor is responsible for making updates to a specific document within $\mathcal{K}$, and a centralized critic uses global feedback to coordinate these updates. The objective is to iteratively improve $\mathcal{K}$ such that the overall accuracy of the RAG system is maximized.
4.3.1 Reward Signal
For a given query $q_{i}$ and the generated answer $o_{i}$, the expert provides feedback $(c_{i},f_{i})$ that includesa ground truth answer $c_{i}$ and qualitative expert feedback $f_{i}$ on any errors. The global reward signal is derived from $c_{i}$ as per the scoring function $s$ (Refer Equation 2).
4.3.2 KB Editing Agent
To effectively incorporate expert feedback, we employ a multi-actor, centralized critic architecture.
Centralized Critic: The centralized critic, denoted as $C$, is responsible for evaluating the overall performance of the RAG system based on the global reward signal $r$ derived from expert feedback. The critic analyzes the feedback received given the current state $s$ of $\mathcal{K}$. The critic’s analysis is then used to provide tailored reflections to each actor, guiding document updates.
The centralized critic aggregates the reward signal across multiple queries to generate a holistic evaluation of $\mathcal{K}$.
$R(s)=\frac{1}{|T|}\sum_{(q_{i},a_{i},c_{i},f_{i})\in T}g(\mathcal{B}(q_{i},\Gamma(q_{i},s)),c_{i})$ (3)
To generate feedback for the documents, the critic needs to take gradient of this reward with respect to the documents. This would give us,
$\partial_{j}=\frac{\partial R(s)}{\partial D_{j}}=\frac{1}{|T|}\sum_{(q_{i},a_{i},c_{i},f_{i})\in T}\frac{\partial}{\partial D_{j}}g(\mathcal{B}(q_{i},\Gamma(q_{i},s)),c_{i})$ (4)
Figure 2 illustrates the environmental interaction of the actor-critic model. Following methodologies in prior works (Pryzant et al., 2023; Juneja et al., 2024; Gupta et al., 2024), we use LLMs to generate an overall text gradient $\nabla$ over each failing example. The critic first identifies and select which documents in $\Gamma(q_{i},s)$ are responsible for any inaccuracies in $o_{i}$. Reflections are then generated for these documents based on the correct answer, expert feedback and the text gradient. However, as shown in Equation 4, we need to aggregate these reflections across all queries. Instead of a simple concatenation, we adopt the clustering approach similar to Juneja et al. (2024), producing generalized reflections that effectively capture the core insights from multiple queries. These aggregated reflections can be effectively considered as the partial textual gradient $\partial$ with respect to the document. These partial gradients are provided as feedback to the document-specific actor $A_{j}$, which then perform the actions to edit the specific documents.
Actors: Each document $D_{i}\in\mathcal{K}$ is managed by a distinct actor, $A_{i}$, which is modeled as a ReACT agent Yao et al. (2023) responsible for making structured edits to its document. Each actor operates independently, receiving reflections from the centralized critic on how to modify the content of $D_{i}=[c_{ij}]$.The actors need to only update these chunks as needed. The set of possible actions includes:
  • $\mathsf{EditChunk}$: The action is defined as $\mathsf{EditChunk}(j,t_{j})$, where $j$ indicates which chunk $c_{ij}$ of $D_{i}$ to modify, and $t_{j}$ is the updated content for the chunk.
  • $\mathsf{AddChunk}$: The action is defined as $\mathsf{AddChunk}(n_{j},t_{j})$, where $n_{j}$ indicates the name of the new chunk, and $t_{j}$ is the content for the chunk.
  • $\mathsf{DeleteChunk}$: The action is defined as $\mathsf{DeleteChunk}(j)$, where $j$ specifies which chunk $c_{ij}$ of $D_{i}$ to remove.
This parameterized action space allows the actors to perform precise edits within the document, ensuring that the refinement process is both flexible and context-specific. Each actor leverages its local state $s_{i}$ and the document-specific feedback from the critic to produce a sequence of structured edits, ensuring that modifications are consistent and contribute towards enhancing the document’s relevance and completeness.
However, there might be many ways to edit a KB and we may need to have some desirable characteristics for the edited KB. In the next section, we discuss what those desirable characteristics could be and how we might measure them.
5 Evaluating Knowledge Base Editing Quality
A Knowledge Base should be complete with respect to a task - it should contain all the information necessary to assist the RAG system to solve the task at hand. Given the open-ended nature of tasks that typical RAG agents are designed for, it is hard to quantify a closed-form metric of completeness. That said, an ideal Knowledge Base editing system should at least be able to incorporate as much external feedback as possible.
Further, It will be extremely undesirable for any Knowledge Base to only help the RAG system for a small subset of tasks. Given the tendencies for data-driven techniques to over-fit on the train-set distribution, it is important that knowledge base edits are generalizable to unseen examples.
Lastly, given the semantic and textual nature of the Knowledge Base, it is important that the documents in the Knowledge base are coherent and consistent throughout. This not only makes the document interpretable for human consumption, it also help reduce in-context noise during LLM inteference, which has been shown to affect LLM performance (Liu et al., 2024).
6 Experimental Setup
6.1 Baseline
While there has been a rich body of works in the area of knowledge editing and prompt optimization, to the best of our knowledge, STACKFEED is the first work targeting the feedback-driven textual Knowledge Base Editing problem. Therefore, to perform a holistic evaluation of STACKFEED we implement - PromptAgent-E, an extension of PromptAgent   Wang et al. (2023) for the KB editing task. PromptAgent  formulates prompt optimization as a strategic planning problem using Monte Carlo Tree Search (MCTS). At a high-level our baseline approach, PromptAgent -E creates separate PromptAgent -style agents to optimize specific document in the KB. To minimize spurious edits in the Knowledge Base, we restrict PromptAgent -E to only optimize documents that were part of the retrievals for more than 2 training sample. After identifying the best nodes for each of the document-wise runs, we put them back in the knowledge base to generate the new version of the KB. In contrast to STACKFEED, PromptAgent -E can be seen as a collection of document-wise Independent Actor-Critic models  (Foerster et al., 2017). We present in-depth comparisons between PromptAgent -E and STACKFEED in Section 7
6.2 Datasets
Knowledge Base Editing can be useful for scenarios where the KB is 1. Incomplete, or 2. Incorrect . We evaluate STACKFEED  on 5 datasets spanning these different settings.
6.2.1 Incomplete Knowledge Base

### v2

@@ 4.1 Knowledge base editing as state search
We model knowledge base (KB) editing as a state optimization problem over the configuration of documents in the KB. Given a query and retrieved evidence, a language model generates a response. When errors arise due to incomplete or inaccurate evidence, we update the KB such that future responses are more accurate.
In this formulation, each KB state reflects a specific configuration of document contents, and actions correspond to edits applied to these documents. A transition function updates the KB by applying these edits, and a reward function evaluates the quality of the resulting KB state based on how well it supports correct and complete responses across a set of queries.
Our objective is to find the optimal KB state that maximizes this reward. This enables a feedback-driven editing process where the system learns to apply targeted modifications to improve RAG performance in a data-driven and interpretable manner. We define the action space, search space and the optimization objective more formally in the appendix section A.1.
4.2 Knowledge Base Editing Agent
We define KB editing agent that operates on a reward signal as a model’s performance over a batch of queries for the given knowledge base in a RAG system.
Centralized Critic: The centralized critic $C$ evaluates the RAG system’s performance by analyzing expert feedback and the current knowledge base state. When errors occur, the critic identifies which specific documents caused the problems and generates targeted feedback for improvement.
The critic examines each failed query and its corresponding expert feedback, first reflecting upon it to fully understand the error and then identifying which documents are responsible for the error. Following established methods from prior work (Pryzant et al., 2023; Juneja et al., 2024; Gupta et al., 2024), rather than simply listing all issues in each retrieved document, the critic clusters similar problems together to identify common patterns and generate more generalizable insights.
These aggregated reflections are analogs to partial gradients $\partial_{j}$ for each document that guide each document-specific actor $A_{j}$ on improving their assigned documents. This approach ensures document updates address systematic issues rather than isolated errors, leading to more effective knowledge base refinement. By analyzing failures across multiple queries and clustering similar issues, the critic provides more strategic guidance than treating each error in isolation.
Pony SciPy Tensorflow CLARK-news
Model Method Acc $\sigma$ Acc $\sigma$ Acc $\sigma$ Acc $\sigma$
GPT-4Turbo Base KB 29.99 1.57 52.04 0.00 28.88 2.18 26.27 1.20
PromptAgent-E 32.22 1.57 53.40 3.12 47.77 3.57 28.80 2.39
STACKFEED 37.04 1.28 59.38 1.22 53.84 3.11 37.28 1.69
GPT-4o Base KB 31.41 1.28 54.13 1.22 31.75 2.91 28.80 1.69
PromptAgent-E 34.21 1.49 55.27 3.05 49.03 3.62 30.01 2.41
STACKFEED 42.32 2.11 61.60 2.43 55.32 2.18 40.40 1.63
GPT-4.1 Base KB 35.40 2.52 53.40 2.43 34.60 3.11 30.89 1.20
PromptAgent-E 36.10 1.73 56.02 3.28 50.27 3.44 31.33 2.19
STACKFEED 45.62 3.67 60.83 2.84 57.61 2.18 43.03 2.14
Table 1: Correctness performance comparison between STACKFEED and baseline method across multiple datasets, reported as accuracy percentages (higher is better). Best results for each model and dataset are highlighted in bold.
Actors: Each document $D_{i}\in\mathcal{K}$ is managed by a distinct actor, $A_{i}$, which is modeled as a ReACT agent Yao et al. (2023a) responsible for making structured edits to its document. Each actor operates independently, receiving reflections from the centralized critic on how to modify the content of $D_{i}=[c_{ij}]$. The actors need to only update these chunks as needed. Each actor is provided with a set of parametrized actions to perform precise edits to the document chunks, allowing for flexible and context-specific edits. The set of possible actions includes:
  • Edit Chunk: Modifies an existing chunk within a document by replacing content with updated text.
  • Add Chunk: Creates a new chunk with specified content and adds it to the document.
  • Delete Chunk: Removes an existing chunk from the document entirely.
5 Experimental Setup
5.1 Baseline
While there has been a rich body of works in the area of prompt optimization, to the best of our knowledge, STACKFEED is the first work targeting the feedback-driven textual Knowledge Base Editing problem. Therefore, to perform a holistic evaluation of STACKFEED we implement - PromptAgent-E, an extension of PromptAgent Wang et al. (2023) for the KB editing task. PromptAgent formulates prompt optimization as a strategic planning problem using Monte Carlo Tree Search (MCTS). We have described our implementation on top of PromptAgent in appendix section A.4
5.2 Datasets
Knowledge Base Editing can be useful for scenarios where the KB is either incomplete or incorrect. We evaluate on EVOR Su et al. (2024) which is a dataset of documentation for programming language Pony which can be incomplete in details along with natural language to code questions in them. Similarly, it has two more datasets about custom versions of SciPy and Tensorflow with the original documentation of these libraries which must be adapted for these custom versions. We also use the also CLARK-News dataset Li et al. (2025a) which is a natural language dataset of news articles of outdated factual information. We describe each dataset in detail in the appendix in A.5.
5.3 System Configurations
For our experiments, we set a maximum search depth of $3$, an expansion width of $3$, and a maximum of $5$ iterations. The UCT algorithm with an exploration constant of $2.5$ is used for expansion nodes. The parameters are chosen to balance between effective exploration and computational cost.
We set up a generic RAG system that uses an embedding similarity for semantic retrieval. Additionally, in line with prior works like (Zhang et al., 2023) for coding-related tasks, we use an iterative retrieval setup wherein we first generate a code using naive retrieval and then query the database again with both the question and the generated code to improve the quality of retrieval before generating the final result. We use OpenAI-Text-Embedding-3-Large as the embedding model and use cosine similarity as a metric of embedding match for ranking.
5.4 Metrics
Correctness We evaluate the correctness of the KB by evaluating it on a test set queries on respective tasks. This separation of the test-train set of queries reduces the risk of contamination of the examples and falsified improvements in performance. We also define two metrics completeness and coherence to understand the quality of the edits made by STACKFEED.
Completeness A knowledge base should be complete with respect to the task, that means it should contain all the information necessary to assist RAG system for task at hand. Given the open-ended nature of tasks that typical RAG agents are designed for, it is hard to quantify a closed-form metric of completeness. However, an ideal KB editing system should at least be able to incorporate external feedback well. To evaluate this we use the precision train set to estimate the degree of expert feedback incorporated in the learned KB.
Coherence Given the semantic and textual nature of the Knowledge Base, it is important that the documents in the Knowledge base are coherent and consistent even after editing. This not only makes the document interpretable for human consumption, it also help reduce in-context noise during LLM inference, which has been shown to affect LLM performance (Liu et al., 2024). To quantify the degree of coherence of the KB, we first calculate coherence scores for each edited document using G-Eval (Liu et al., 2023). We use the G-eval score to gauge the coherence of an edit made to a KB document to the document itself. And the mean of this document-level coherence over all the documents is defined as the coherence of an edited KB.
[…]
Dataset Pony SciPy Tensorflow CLARK-news Pony SciPy Tensorflow CLARK-news
STACKFEED GPT-4Turbo 9.68 31.38 44.44 13.79 4.6 4.30 4.0 1
STACKFEED GPT-4o 11.38 36.67 50.12 15.41 4.67 4.67 4.0 2.33
STACKFEED GPT-4.1 13.45 41.46 52.24 18.62 4.6 4.00 3.67 1
6 Results and Analysis
We evaluate STACKFEED on three different OpenAI models GPT-4Turbo, GPT-4o and GPT-4.1111https://platform.openai.com/docs/models on three different configurations. Firstly, evaluating the performance on the Base KB, this is the initial state of the KB $s_{0}$ without any edits. We then make a series of edits on $s_{0}$ using PromptAgent-E and a series of edits by STACKFEED on KB state $s_{0}$.
We report our main results in table 1. We observe the performance of the RAG system constantly improve with edits done by STACKFEEDfor all the three models.
Quality of edits As seen in Table 2, STACKFEED produces edits with a coherence score of 4 or higher. For KBs that need long-term maintenance (such as language and code documentation as seen in the Evor datasets), STACKFEED makes more coherent edits compared to the baseline. This is especially true for long documents, as seen in the EVOR Pony dataset. We also note that the edits made by PromptAgent-E were made in incorrect documents leading to more noisy generations.
We also observe that PromptAgent-E added irrelevant section in example A.6 on the lineSearch and norm_ppf functions in a document about sparse matrices. These edits were made because the document was retrieved for questions which had errors regarding these functions. These edits are irrelevant to the document and showcase reason for failure in PromptAgent-E is to make documents less coherent.
On the other hand, STACKFEED makes more relevant edits to the document which are contained to the context of the document. This shows how STACKFEED is able to maintain the coherence of the document in its edits. We demonstrate an complete example of edits made in appendix A.6
For a news-article-like dataset like CLARK-news with factual edits. Incoherency is naturally induced when the facts of the article change. In this dataset, coherence is sacrificed in bringing the facts of the article up to date, which is required to improve accuracy.
We also evaluate performance on EVOR-Pony dataset using just a single STACKFEED edit and greedy search in table 3. In greedy search, we greedily pick the most rewarding node at a particular depth. We observe even though the completeness of the document is quite similar the edits are much less coherent and the generalize less than MCTS based search.
Single Edit Greedy Search MCTS
Correctness 36.14 41.34 45.62
Completeness 9.68 13.96 13.45
Coherence 4 3.34 4.6
Table 3: Comparison between single edit, greedy search and MCTS, on EVOR Pony dataset with GPT-4.1. We used max width=3 and max depth=5.
6.1 Generalization to Repository Level Migration
Migration Efficacy (in %)
Model Method $\eta_{\text{minimal}}$ $\eta_{\text{maximal}}$
GPT-4o Base KB 43.67% 18.12%
STACKFEED 46.14% 23.02%
GPT-4.1 Base KB 58.81% 26.50%
STACKFEED 61.71% 28.16%
Table 4: Test set performance of STACKFEED and Base KB on Migration Bench
To further test the efficacy of STACKFEED , we apply STACKFEED to a more complex real-world agentic scenario, namely repository-level code editing. We choose to work with the (Liu et al., 2025a) benchmark, which consists of migration of 5,102 real repositories from Java 8 to Java 21 along with a comprehensive evaluation harness for any system attempting this task. This benchmark presents unique challenges including a substantially larger action space requiring navigation of entire repository structures, more delayed feedback signals determined by compilation and test execution, and specialized knowledge requirements for Java version-specific APIs and migration patterns.
For the purpose of this study, we extend the SDFeedback agent introduced in (Liu et al., 2025a) to incorporate KB-guided editing, where the agent queries the knowledge base before making each edit decision. For the KB, we manually create a small knowledge base consisting of basic information relevant to Java migration, hereby referred to as the BaseKB for this setting. The BaseKB contains 47 entries covering common migration patterns such as API replacements, deprecated method alternatives, and frequently encountered compilation errors with their resolutions. We deliberately keep the BaseKB small and somewhat incomplete to better demonstrate STACKFEED’s ability to identify and fill knowledge gaps.
We use the Selected subset of the benchmark as our test set, and we randomly sample 100 examples from the remaining Full subset of the benchmark, dividing them equally to create training and validation sets. We run STACKFEED with the same MCTS settings as the other datasets. We run this experiment with two models, GPT-4o and GPT-4.1. We use the metrics introduced in (Liu et al., 2025a) to represent the success rate of the migration, namely minimal migration ($\eta_{minimal}$), which checks whether the Java version is correctly updated and whether the migrated repository passes all test cases, and maximal migration ($\eta_{maximal}$), which adds the additional criterion of all dependencies needing to be updated to their latest versions. The complete details of the experiment, including exact modifications made to SDFeedback and the manually created knowledge base, are present in Appendix Section A.2.
Table 4 shows the migration success rates for the Base knowledge base and the knowledge base after STACKFEED was applied. We observe approximately 3% absolute improvement in $\eta_{minimal}$ for both models over the Base knowledge base, showcasing the effectiveness of the edits in addressing both functional correctness and dependency management. These gains are particularly notable given the already strong baseline performance of the SDFeedback agent.
Figure 3 shows that the edits made to the base knowledge base consist primarily of addition of new error types and patterns not present in BaseKB (15 new entries), expansion of existing entries with more detailed resolution steps including specific code examples (23 entries expanded), and corrections to incomplete guidance (9 entries corrected). This demonstrates how STACKFEED adds missing information to the knowledge base while also expanding upon and correcting already existing information, allowing the agent to make better edits. Edits on a section of the knowledge base can be seen in Figure 3.
7 Conclusion
We introduced STACKFEED, a novel framework for refining Knowledge Bases (knowledge bases) in Retrieval-Augmented Generation (RAG) systems using a multi-actor, centralized critic architecture. STACKFEED enables efficient knowledge base updates without retraining or altering model parameters by leveraging feedback-driven structured edits and textual gradients.
Our approach achieved superior performance in preserving knowledge base in terms of coherence, consistency, and completeness, resulting in enhanced performance of RAG system.
Broader Impact
The deployment of Retrieval-Augmented Generation (RAG) systems in real-world applications such as AI-powered developer assistants, enterprise chatbots, and domain-specific information retrieval relies heavily on the correctness and reliability of the underlying knowledge bases (KBs). However, maintaining these KBs is a persistent bottleneck due to frequent changes in domain-specific knowledge and the lack of automated mechanisms for continuous KB refinement. Our proposed system, STACKFEED, addresses this challenge through a feedback-driven framework for automatic knowledge base editing that learns from expert or oracle feedback in real-time deployments.
Our design introduces a modular, actor-critic architecture that can be integrated into existing pipelines with minimal engineering overhead. By defining clear KB quality metrics—correctness, coherence, and completeness. Our system provides actionable insights for both developers and auditors. This supports not only continuous improvement of deployed systems, but also regulatory compliance and human-in-the-loop oversight in high-stakes domains like healthcare, finance, and legal automation. By automating feedback incorporation into KBs, we reduce human maintenance cost, lower response errors in production RAG systems, and promote safer, more trustworthy deployments.
In industrial RAG based applications, post-deployment error-correction and maintenance is done through improving the quality of the retrieval system. We introduce another axis by enabling the optimization of the Knowledge base itself. It also paves the way for joint optimization of both the knowledge base and the retrieval mechanism, offering a more holistic and scalable solution to long-term system maintenance.
We hope this work paves the way for future industry adoption of learning-enabled infrastructure that maintains and improves itself in deployment, and encourages further exploration of editable memory systems as an alternative to end-to-end retraining for knowledge maintenance.
8 Limitations and Future Work
While this work presents a novel framework for feedback-driven knowledge base refinement, several limitations and corresponding avenues for future research should be acknowledged. In particular, one limitation of this work is the decoupling of the knowledge base optimization and retrieval. This work assumes that the retrieval component can correctly identify the right documents to retrieve. Failures originating from faulty retrieval cannot be holistically addressed by this system. A promising avenue for future work can be the joint optimization of both the knowledge base and the retrieval mechanism. Creating a unified framework that can decide whether to fix an error by editing a document or by tuning the retriever could offer a more holistic and scalable solution to long-term system maintenance.
Appendix A Appendix
A.1 Knowledge Base Editing as State Search
@@
Given a query $q_{i}$ and a set of retrieved documents $\Gamma(q_{i},\mathcal{K})$, the LLM $\mathcal{M}$ generates an answer $o_{i}$. When errors arise due to incomplete or incorrect information in the retrieved documents, our goal is to identify the optimal configuration of $\mathcal{K}$ that improves the accuracy of the system’s responses. Thus, we define our state search problem as finding the best state $s^{*}$ of the KB.
@@
Environment: We model the environment simply as a “patch” function, that takes the diff generated by the agent and patches the KB to produce the new state.
@@
$R(s)=\frac{1}{|Q|}\sum_{q_{i}\in Q}g(\mathcal{M}(q_{i},\Gamma(q_{i},s)))$ (2)
$s^{*}=\arg\max_{s\in\mathcal{S}}R(s)$ (3)
@@
The reward function $R(s)$ is derived from the expert feedback on the system’s generated answers and captures improvements in the KB. By optimizing for $s^{*}$, we ensure that the final state of the KB maximizes the overall accuracy and effectiveness of the RAG system, rather than focusing on an intermediate sequence of state transitions.
@@
A.2 Migration Bench
We share the KB used for Java migration
Java KB
## **‘java.lang.UnsupportedClassVersionError‘**
**Symptom:**
Exception in thread "main" java.lang.UnsupportedClassVersionError:
com/example/App has been compiled by a more recent version of the Java Runtime
(class file version 65.0), this version of the Java Runtime only recognizes up to 61.0
**Cause:**
* The code is compiled with Java 21 but run with an older JVM (e.g., Java 17).
* Class file version 65.0 corresponds to Java 21.
**Fix:**
* Ensure the runtime JVM matches the compiler version:
‘‘‘bash
java -version
javac -version
* Upgrade your runtime to Java 21.
## **‘module not found: java.base‘**
**Symptom:**
error: module not found: java.base
**Cause:**
* Misconfigured module path.
* Incorrectly set ‘--release or ‘--module-path flags.
**Fix:**
* Verify JAVA_HOME points to JDK 21.
* Use:
‘‘‘bash
javac --release 21 ...
* Ensure dependencies are on the module path (or use classpath if not modularized).
## **‘invalid source release: 21‘**
**Symptom:**
error: invalid source release: 21
**Cause:**
* Older build tool (Maven, Gradle, Ant) that does not yet support Java 21.
**Fix:**
* Upgrade the build tool version:
* **Maven:** $\geq$$ 3.9.x with maven-compiler-plugin 3.11+
* **Gradle:** $\geq$$ 8.3
* Example Maven pom.xml‘:
‘‘‘xml
<plugin>
<groupId>org.apache.maven.plugins</groupId>
<artifactId>maven-compiler-plugin</artifactId>
<version>3.11.0</version>
<configuration>
<release>21</release>
</configuration>
</plugin>
## **‘Preview feature used without --enable-preview‘**
**Symptom:**
error: patterns in switch are a preview feature and are disabled by default.
**Cause:**
* Code uses **preview features** in Java 21 (e.g., string templates, unnamed classes, pattern matching in switch).
**Fix:**
* Compile and run with preview enabled:
‘‘‘bash
javac --enable-preview --release 21 MyApp.java
java --enable-preview MyApp
* Avoid preview features in production code.
## **Gradle Daemon / Toolchain Errors**
**Symptom:**
Could not target platform: Java SE 21’ using tool chain: JDK 17 (17)’.
**Cause:**
* Gradle is using an older JVM despite Java 21 being installed.
**Fix:**
* Configure toolchain in build.gradle‘:
‘‘‘groovy
java {
toolchain {
languageVersion = JavaLanguageVersion.of(21)
* Ensure JAVA_HOME points to JDK 21.
## **‘cannot find symbol with Standard APIs**
**Symptom:**
error: cannot find symbol
symbol: class SequencedCollection
**Cause:**
* Using new Java 21 APIs (e.g., SequencedCollection‘, VirtualThread‘) but compiling with an older JDK.
**Fix:**
* Compile and run with JDK 21.
* Ensure IDE is configured with Java 21.
## **‘Illegal reflective access warnings**
**Symptom:**
WARNING: An illegal reflective access operation has occurred
**Cause:**
* Libraries using reflection to access internal JDK APIs, stricter in newer Java versions.
**Fix:**
* Update to latest versions of affected libraries.
* If unavoidable, use JVM args (not recommended for long-term):
‘‘‘bash
--add-opens java.base/java.lang=ALL-UNNAMED
Listing 1: MigrationBench KB
Figure 3: The above example showcases the edits made by STACKFEED.① represents a more precise and structured state of information made from edits by STACKFEED. ② showcases a fix that was written more coherently and with added details for the agent by observations made from trajectory.③ showcase added information from train set on resolution.
A.3 Example and Overview
Figure 2 illustrates our technique applied to the Pony (Su et al., 2024), where a knowledge base (KB) for the low-resource programming language Pony supports a natural language-to-code task. Due to Pony’s rarity, language models often generate code that fails to compile. To address this, we use the Pony compiler as an expert to provide feedback in the form of compile errors.
1
2
Centralized Critic Analysis: For all errors, the critic analyzes why the error occurred. For instance, if the apply method in the Array class is partial and may raise an error, the critic suggests adding a ? to handle potential failures. Based on this reflection, the critic identifies which of the retrieved document is relevant for the error and provides a tailored reasoning for it.
3
Per-Document Actor: For each document in the KB, the gradients associated to it are aggregated. This aggregate gradient is used as a signal by the Per-Document Actor, in this case, the actor for document builtin-array.md to make edits to the document.
4
Re-evaluation and MCTS Search: After edits are applied, the KB is re-evaluated, generating new feedback and a reward score. This score guides a Monte Carlo Tree Search (MCTS) to explore different states of the KB, iterating through steps 1- 3 to progressively refine the KB and improve the system’s overall performance.
A.4 PromptAgent-E Baseline
PromptAgent Wang et al. (2023) is a technique developed for optimizing a single prompt. PromptAgent-E extends this approach to knowledge base (KB) optimization by independently optimizing each document within the KB using a separate PromptAgent instance. Unlike STACKFEED, PromptAgent-E operates as a collection of document-wise Independent Actor-Critic models Foerster et al. (2017).
A.4.1 Algorithm Description
The PromptAgent-E algorithm proceeds as follows:
  1. 1.
    Initial Evaluation: Given a training set of queries, we first run the current KB to obtain retrievals, generations, and expert feedback for these generations. The same is done for the validation set.
  2. 2.
    Document-Level Dataset Creation: The training set is then segmented into document-level training sets. Each document-level set comprises all queries (along with their corresponding retrievals, generations, and expert feedback) for which a specific document was retrieved. Similarly, the validation set is also split into document level validation sets.
  3. 3.
    Document Selection for Editing: Given that KBs can be extensive, we restrict editing to documents that were retrieved for at least two queries in the training set. This ensures focusing on more relevant or frequently accessed documents.
  4. 4.
    Independent Optimization: A separate PromptAgent instance is then created and executed for each selected document independently. Each document-level PromptAgent instance only accesses the queries, generations, and feedback pertinent to its assigned document.
  5. 5.
    KB Update: After determining the optimal node (or prompt) for each document, these optimized nodes are integrated back into the KB to form a new, improved version.
A.5 Dataset
Knowledge Base Editing can be useful for scenarios where the KB is
  1. 1.
    Incomplete: the knowledge bases misses some key artifacts responsible for answering the questions. In the Evor-Pony dataset, the documentation used lacks information on various aspects of the language like partial functions etc.
  2. 2.
    Incorrect: the knowledge base in this case consists of some incorrect knowledge.
. We evaluate STACKFEED on 5 datasets spanning these different settings.
A.5.1 Incomplete Knowledge Base